-l1 points SerPSE9 24.XP.006 A solid plastic spherHow do I do this problem?
Answer #1

The equation for the electric field for two different position is

\frac{E_2}{E_1} = \frac{k (Q/r^2)}{k(Qr/a^3)}

\frac{E_2}{E_1} = \frac{a^3}{r^3}

\frac{E_2}{25 kN/C} = \frac{10^3}{14^3}

E_2 = 9.11 kN/C